# The Nuclear Star Cluster of M 74: a fossil record of the very early stages of a star-forming galaxy

### v2

@@ Appendix A Analysis of the original MUSE data cube of the galaxy
We present here the results from the analysis of the original cropped data cube of the central arcmin of M 74 (Sect. 4.1.1). The analysis was performed following the method described in Sect. 4.2. While the goal of the paper is to analyze separately the NSC and the host galaxy, the results for the integrated data cube provide a general context and serve as a reference. In Fig. 8, we show the maps of the mean age, total metallicity ([M/H]) and [Mg/Fe] abundance, from top to bottom, extracted from the original MUSE data cube. This cube is the kinematic-corrected data cube resulting from the preparation described in Sect. 4.1.1, before being decomposed in different data cubes with C2D. A Voronoi binning to a target S/N=100=100 was also used in this case. However, the different preprocessing compared to the host-galaxy data cube resulting from the C2D decomposition (Sect. 4.1.2) leads to a different bin size. Considering that Voronoi bins are, in general, larger in Fig. 8 than in Fig. 6, the maps in the two figures are totally compatible with each other. Similarly to Fig. 6, the age map (top panel of Fig. 8) shows alternatively regions dominated by relatively young (younger than ${\sim}4$ Gyr) and old stars (older than ${\sim}8$ Gyr), roughly following the pattern of the spiral arms. [M/H] shows regions of higher and lower values. As in Fig. 6, there are regions with low metallicity in the outskirts of the mapped region, while we find supersolar metallicities within the central ${\sim}500$ pc. [Mg/Fe] abundances are in general anti-correlated with [M/H], with values relatively close to solar in the central ${\sim}500$ pc. Zooming into the very center of the middle panel in Fig. 8, a clear [M/H] drop is seen in the PSF size. However, no significant age and [Mg/Fe] differences are seen in the very center of the top and bottom panels with this spatial binning level.
Figure 8: Stellar population maps of the central arcmin of M 74, extracted from the integrated data cube. From top to bottom: mean age, total metallicity [M/H] and [Mg/Fe] abundance. Masked regions are depicted in white and isophotal contours in black. The physical scale is given as a reference in the top $X$ axes.

### v3

@@ Appendix A Analysis of the original MUSE data cube of the galaxy
We present here the results from the analysis of the kinematic-corrected original data cube of M 74 (Sect. 4.1.1). A Voronoi binning to a target S/N=100=100 was also used in this case. However, the different preprocessing compared to the host-galaxy data cube resulting from the C2D decomposition (Sect. 4.1.2) leads to a different bin size compared to the host-galaxy data cube (Fig. 6). The analysis was performed following the method described in Sect. 4.2. While the goal of the paper is to analyze the NSC and the host galaxy separately, the results for the integrated data cube provide a general context and serve as a reference. In Fig. 8, we show the maps of the mean age, total metallicity ([M/H]) and [Mg/Fe] abundance, from top to bottom, extracted from the original MUSE data cube. Considering that Voronoi bins are, in general, larger in Fig. 8 than in Fig. 6, the maps in the two figures are totally compatible with each other. Similarly to Fig. 6, the age map (top panel of Fig. 8) shows alternatively regions dominated by relatively young (younger than ${\sim}4$ Gyr) and old stars (older than ${\sim}8$ Gyr), roughly following the pattern of the spiral arms. [M/H] shows regions of higher and lower values. As in Fig. 6, there are regions with low metallicity in the outskirts of the mapped region, while we find supersolar metallicities within the central ${\sim}500$ pc. [Mg/Fe] abundances are in general anti-correlated with [M/H], with values relatively close to solar in the central ${\sim}500$ pc. Zooming into the very center of the middle panel in Fig. 8, a clear [M/H] drop is seen in the PSF size. However, no significant age and [Mg/Fe] differences are seen in the very center of the top and bottom panels with this spatial binning level.
Figure 8: Stellar population maps of the central arcmin of M 74, extracted from the integrated (kinematic-corrected) data cube. From top to bottom: mean age, total metallicity [M/H] and [Mg/Fe] abundance. Masked regions are depicted in white and isophotal contours in black. The physical scale is given as a reference on the top $X$ axis.